Web Cites Research projects a rate of return of 21 percent on new projects. Management plans to plow back 45 percent of all earnings into the firm. Earnings this year will be $4.1 per share, and investors expect a 10 percent rate of return on the stock.

What is the sustainable growth rate?

What is the stock price?

What is the present value of growth opportunities?

What is the P/E ratio?

What would P/E ratio be if the firm paid out all earnings as dividends?
Answer #1

a)
Sustainable growth rate = ROE * retention ratio
= 21% * 45%
= 9.45%

b)
D1 = EPS1 * (1 - ratention ratio) = $4.1 * (1 - 0.45) = $2.255

P0 = D1 / (Ke - g)
= $2.255 / (0.10 - 0.0945)
= $2.255 / 0.0055
= $410

Stock price = $410

c)
Present value of growth opportunity = P0 - (EPS / Ke)
= $410 - ($4.1 / 0.10)
= $410 - $41
= $369

d)
P/E ratio = Market price / EPS
= $410 / $4.1
= 100

e)
P0 = D1 / (Ke - g)
= $4.1 / (0.10 - 0)
= $41

P/E ratio = $41 / $4.1 = 10